Hi,
Would you be able to help me rearrange this equation to make r the subject.
260= 450/(1+ 2*(e^-r*t))
Thankyou

Answer by Alan3354(69443)

Would you be able to help me rearrange this equation to make r the subject.
260= 450/(1+ 2*(e^-r*t))
------------
Yes, I would.
1+ 2*(e^-r*t) = 450/260
Your move.